OpenStudy (myrraarrym):

HELP PLEASE !!! A number cube has the numbers 1, 2, 5, 5, 6, and 6 on its faces. Match the probability with the outcome.

OpenStudy (myrraarrym):

help please!!!!!!!!!

jimthompson5910 (jim_thompson5910):

hint: P(4) means "The probability of rolling a 4". None of the faces listed have "4" on them, so it's impossible to roll a 4. This means that P(4) = 0 whenever the probability is 0, it means the event can never happen

OpenStudy (sshayer):

\[P(5)=\frac{ 2 }{ 6 }\] \[P(6)=\frac{ 2 }{ 6 }\] \[P(5 U 6)=P(5)+P(6)=?\]